OpenStudy (anonymous):

benjamin adds 4 new songs too his mp3 player each week after 7 weeks he has 43 songs which equation in slope intercept form represents the total number of songs he has on his mp3 player after x weeks

OpenStudy (anonymous):

43=4(7)+c c=15 y=mx+c y=4x+15